Problem 10-2A Depreciation methods LO P1

A machine costing $216,200 with a four-year life and an estimated $19,000 salvage value is installed in Luther Company’s factory on January 1. The factory manager estimates the machine will produce 493,000 units of product during its life. It actually produces the following units: 122,000 in 1st year, 122,600 in 2nd year, 119,900 in 3rd year, 138,500 in 4th year. The total number of units produced by the end of year 4 exceeds the original estimate—this difference was not predicted. (The machine must not be depreciated below its estimated salvage value.)

  Required: Compute depreciation for each year (and total depreciation of all years combined) for the machine under each depreciation method. (Round your per unit depreciation to 2 decimal places. Round your answers to the nearest whole dollar.)
Answer #1

Depreciation cost per unit = (cost - salvage value) / Total estimated machine hours

= (216,200 - 19,000) / 493,000

= 0.4 per unit

Year Depreciation expense
1 48,800 (122,000*0.4)
2 49,040 (122,600*0.4)
3 47,960 (119,900*0.4)
4 51,400 *

* Depreciation expense in year 4 = 138,500*0.4 = 55,400

Depreciation expense in year 4 is limited to 55,400 - (10,000*0.4) = 51,400